### v1

[…]
The dependence of (980) production on the final-state charged-particle multiplicity is reported for proton–proton (pp) collisions at the centre-of-mass energy, 13 TeV. The production of (980) is measured with the ALICE detector via the decay channel in a midrapidity region of 0.5. The evolution of the integrated yields and mean transverse momentum of f0(980) as a function of charged-particle multiplicity measured in pp at 13 TeV follows the trends observed in pp at 5.02 TeV and in proton–lead (p–Pb) collisions at 5.02 TeV. Particle yield ratios of (980) to and (892)0 are found to decrease with increasing charged-particle multiplicity. These particle ratios are compared with calculations from the canonical statistical thermal model as a function of charged-particle multiplicity. The thermal model calculations provide a better description of the decreasing trend of particle ratios when no strange or antistrange quark composition for f0(980) is assumed, which suggests that the tetraquark interpretation of the f0(980) is disfavored.
1 Introduction
Quantum chromodynamics (QCD) is the theory of the strong interaction. It describes the interactions between coloured quarks and gluons, and the formation of colour-neutral hadronic states. In addition to the conventional meson (q) and baryon (qqq) states described by the quark model [1], exotic states such as tetraquarks [2, 3, 4], pentaquarks [2, 5], hexaquarks [6], and other compositions of a higher number of constituent quarks [7] are allowed, which are of interest in theoretical and experimental research [8].
Light scalar mesons, one of the most frequently revisited hadron classes, consist of light-flavoured quarks and are characterized by quantum numbers and masses below 2 GeV/. The study of light scalar mesons, including their internal structure, properties such as decay width, and formation mechanisms, is an actively explored research area [9, 10, 11]. The is a light scalar meson and has a large mass ( MeV/) and a relatively small particle width ( 40–70 MeV/) [12, 13]. Given these characteristics, the can currently be experimentally identified compared to other light scalar states, such as and , where the threshold mass is close to the invariant mass. Theoretical interpretations of the elusive internal structure of this particle suggest that can either be a conventional meson () [14], a compact tetraquark [15], or a molecule [16]. These competing scenarios remain the subject of intense theoretical debate [12].
QCD predicts the formation of a deconfined and strongly interacting matter, called the quark–gluon plasma (QGP), under conditions of high temperature and high energy density [17, 18, 19, 20]. As the system cools down, the near-perfect fluidity of QGP collapses, confining quarks and gluons into hadrons upon crossing its critical temperature. As the system cools below the critical temperature, all inelastic interactions stop, confining partons into hadrons, where the relative particle composition is expected to be fixed at the chemical freeze-out stage. In heavy-ion collisions, the thermally produced hadron gas can be modelled according to a grand canonical ensemble of particle species, considering the conservation of baryon number, electric charge, strangeness, and other quantum numbers on average. Statistical hadronisation modelling of particle production can be extended to pp and p–Pb collisions by employing a canonical approach. An implementation of the statistical model is the canonical statistical model with undersaturation of strangeness, CSM [21]. The CSM accounts for the system-size-dependent hadrochemistry at vanishing baryon density with local conservation of electric charge, baryon density, as well as strangeness. It can, therefore, predict the relative particle composition while suppressing the production of strangeness in small collision systems. Although CSM does not account for the hadronic phase or the rescattering of decay products from short-lived resonances, it remains a powerful tool for investigating strangeness production. In particular, such an approach can provide insight into whether certain hadrons include strange quarks. Due to the exotic nature of and its production mechanisms, it has not yet been implemented in most QCD-inspired general-purpose Monte Carlo models utilised in the field. Therefore, the CSM offers a suitable framework for studying the strangeness-related properties of the . For instance, the enhancement of ratio with increasing multiplicity can indicate the presence of strange quarks inside the . On the other hand, the absence of late hadronic interactions in the statistical hadronisation model results in the overestimation of resonance yields, especially in large collision systems. Therefore, the model overestimates the observed (980)/ ratio. One open point, which is relevant to a precise data-to-model comparison [22], is the knowledge of the (980) branching ratios (B.R.). Only few measurements of the B.R. of the decay channel [23, 24, 25] are available. The CELO [23] and BaBar [24] measurements of the ratio of B.R. are compatible with zero within 3 units of their uncertainties, where the zero value of the ratio indicates no decay. Such a measurement of the B.R. is relevant because it directly affects the corrected yields. However, after chemical freeze-out, the system still evolves through a hadronic phase, during which elastic interactions and resonance dynamics play an important role.
In the hadronic phase, the system is a dense gas of hadrons and resonances. In particular, strongly decaying states have short lifetimes ( 1–10), often comparable to the duration of the hadronic phase ( 1–10). As a result, these particles can decay within the medium, and their decay products may undergo rescattering [26] with surrounding hadrons. At the same time, resonance regeneration [26] can occur from interactions between particle pairs in the hadronic phase. In this paper, we refer to these two processes collectively as hadronic interactions, which are believed to modify the measurable yields of resonances. Short-lived resonances, such as (770)0 [27], (892)0 [28, 29, 30], K∗(892)± [31], [32, 33], (1520) [34], and [13], serve as important probes of the late hadronic phase. Their yields can be compared to those of long-lived or ground-state particles to study the effects of hadronic interactions. The ratios of and are particularly illustrative. Notably, extracting ratios of particles with the same strangeness content mitigates potential strangeness enhancement effects, thus permitting direct interpretation of these modifications. Recently, system-size dependent modifications of particle yields have been observed in high-multiplicity pp and p–Pb collisions and have been interpreted as a result of hadronic interaction effects that occur in these small collision systems [29, 28]. The measurement of the modification of the yield may contribute to further understanding of the late hadronic phase owing to its short lifetime ( 5–10) [12].
This paper reports measurements of production in pp collisions at the centre-of-mass energy of 13 TeV. The has been measured at midrapidity ( 0.5) in the transverse momentum () range of 0 13 GeV/ and for different final-state charged-particle multiplicity classes. Section 2 describes the experimental setup. In Sec. 3, the reconstruction of and the relative corrections are detailed. The systematic uncertainties in the measurement are described in Sec. 4. Section 5 presents the results, including spectra, -integrated yields, mean , and the particle yield ratios. Conclusions are drawn in Sec. 6.
2 Experimental setup
The data sample used in the analysis was collected with the ALICE detector during the LHC Run 2 (2015–2018) of pp collisions at = 13 TeV with a minimum bias (MB) trigger. Details of the ALICE apparatus during the LHC Run 2 are described in Refs [35, 20]. The following detectors are used in the analysis: the V0 [36, 37], the Inner Tracking System (ITS) [38], the Time Projection Chamber (TPC) [39], and the Time-Of-Flight (TOF) [40].
The V0 detector consists of two arrays (V0A and V0C) of 32 scintillators located on both sides of the interaction point, covering the full azimuthal angle within the pseudorapidity intervals and , respectively. The MB pp collisions are selected online by requiring a signal in both V0A and V0C arrays in coincidence with a LHC bunch crossing. The total charge deposited in the V0A and V0C, which is called V0M, is utilised to determine the multiplicity classes. The multiplicity classes are defined based on the V0M amplitudes, which are proportional to the multiplicity of charged particles in the forward and backward rapidity regions. The event class of INEL 0 is defined as at least one charged track at midrapidity. The MB sample utilised in the present analysis corresponds to an integrated luminosity of = 19 [41].
The ITS is composed of six cylindrical layers of silicon detectors, which are arranged in the radial direction from the beam pipe. The innermost two layers of the ITS, the Silicon Pixel Detector (SPD) [42], provide high-precision points in the vicinity of the primary vertex (PV) of the collision. The PV position along the beam direction () is required to be in cm from the nominal interaction point ( ). Pileup events are filtered by rejecting events with multiple reconstructed vertices with the additional requirement that the distance between the PV and any additional reconstructed vertex is larger than cm. In addition, a mismatch between the number of track candidates in the ITS and the clusters in the SPD is utilised to further suppress the pileup events [43]. After these selections, the probability of pileup events is expected to be about 0.1% in the MB sample [44]. Charged particles are reconstructed down to GeV/ in the pseudorapidity range over the full azimuth with the TPC and the ITS detectors. These detectors are located inside a large solenoidal magnet, providing a uniform magnetic field of T directed along the beam axis. Particle identification (PID) is performed with the TPC and TOF. The TPC measures specific ionisation energy loss of charged tracks to identify particle species. The TOF is used for additional PID by measuring the flight time of charged particles from the PV to the TOF.
3 Data analysis
The resonance is reconstructed via the decay channel , with the reported B.R. of ()% [25]. The candidates are identified through pairs of charged tracks reconstructed using the ITS and TPC. The selection criteria for the reconstructed tracks include 0.15 GeV/ and 0.8, to ensure uniform detector acceptance. Further details on the quality of selection criteria used in this analysis are available in [45]. To guarantee good momentum resolution, the tracks must traverse at least 70 readout pad rows in the TPC (out of a maximum of 159) and have a minimum of two hits in the ITS (out of a maximum of 6), including at least one in the SPD. To minimise the contamination from secondary charged particles originating from weakly decaying hadrons or interactions with the detector material, selection criteria are applied to the distance of closest approach (DCA) to the PV. The transverse component, , is subject to a -dependent cut, while the longitudinal component, , has a fixed threshold. Specifically, must be less than 2 cm, and must be less than cm, where is in GeV/.
The identification of charged pions is performed by using the combined information of the TPC and TOF [46]. Inside the TPC, the ionisation energy loss is required to be within two standard deviations of the expected value from a Bethe–Bloch parameterisation for pions. In the TOF, the measured flight time should also fall within three standard deviations of the expected flight time for a pion. If the TOF signal is unavailable, pion identification relies solely on the energy loss information provided by the TPC.
The signal is extracted through an invariant mass analysis by pairing two oppositely charged pions from the same event within the pair rapidity range 0.5. The combinatorial background, arising from random combinations of uncorrelated pion pairs, is subtracted using the like-sign method [27, 47]. The like-sign background is generated by taking the geometric average of and distributions from the same events, which is equal to 2. After subtracting the like-sign background from the signal or unlike pairs distribution, the resulting distribution reveals peaks indicative of resonances decaying to . Figure 1 shows the like-sign-subtracted invariant mass () distributions for 1.0 1.5 GeV/ (4.0 5.0 GeV/) in MB events on the left (right) panel.
Since (770)0 and (1270) both decay dominantly to and have broad widths, the candidates include contributions from these two resonances. Furthermore, a residual background (), which is mainly arising from misidentified particles and mini-jets, is present. To disentangle the contributions of these backgrounds and the three resonances, the measured invariant mass distribution is modelled with a superposition of several functions. Each resonance contribution is described with a relativistic Breit-Wigner function (rBW) [27, 48]. It is important to note that the detector resolution of (1 MeV) gives a negligible contribution to the widths of these broad resonances [29]. The rBW can be expressed as
where is the relativistic particle width and can be expressed as
Here, , , and denote the amplitude, the rest mass, and the width of the resonance, respectively. The spin of the resonance is represented as , and the charged pion mass (), is 139.6 MeV/ [12]. The spins for the , (770)0, and (1270) resonances are 0, 1, and 2, respectively. In order to correct for the phase space effects, each resonance rBW is multiplied by the phase space factor () [27], which can be expressed as
where represents the transverse momentum of the pair and is set to the median for each interval, and , the kinetic freeze-out temperature, is set to 160 MeV for all defined multiplicity classes [27], and is varied to estimate the corresponding systematic uncertainties. The can be determined based on a Blast-wave analysis and is observed to vary with the multiplicity class [45]. The fit function for the residual background () is modelled using a Maxwell-Boltzmann-like distribution [49], which can be expressed as
where, , , , and are free parameters.
The total fit function is composed of three resonance rBWs and the . The fit function has nine free parameters, including three for the resonance (mass, width, and amplitude), two amplitudes for the (770)0 and f2(1270) resonances, and four parameters for . The width of the resonance is not yet precisely determined by measurements and is, therefore, left as a free parameter in the fit, with a range of 10 100 MeV/ [12]. The masses and widths of (770)0 and (1270) are fixed to their world-average values from Ref. [12], namely 766.5 MeV/, 149.1 MeV/, 1,275.5 MeV/, and 186.7 MeV/.
Due to the large number of free parameters in the fit function, the fitting procedure is divided into three distinct steps, designed to prevent the convergence of parameter values at local minima. At the beginning, an initial value for the width is determined using the MB sample, employing the same binning to minimise the impact of statistical fluctuations. At this stage, all nine parameters are allowed to vary freely. The subsequent step focuses on refining the parameters of , while the width is fixed at the value derived from the first step. The final step involves fixing the parameters as determined earlier and permitting the width to adjust within a specified range of 10 100 MeV/. Here, the fit allows the amplitudes of the three resonances and the mass to vary freely, with the fit range set between 0.8 1.76 GeV/. Notably, the extracted width of the resonance varies from 40 to 70 MeV/ across different and multiplicity intervals used in this analysis.
The raw yields of () for each interval are calculated by integrating the rBW function. These yields are subsequently corrected for the reconstruction efficiency () and normalised for the number of selected pp collisions (), the width of the (), rapidity interval ( 1), and the B.R. [25]. The fully corrected yield is given by the following equation:
The yield of the resonances is corrected and normalised to the number of events in each multiplicity class, represented by . This is done using the factors . The trigger efficiency and vertex reconstruction efficiency depend on the multiplicity class, increasing from 0.85 to 1 and 0.97 to 1, respectively, as the multiplicity increases. The coefficients for the acceptance () and the efficiency () of the tracking and PID for pion pairs are estimated using a detailed simulation of the ALICE detector response. The pp collisions are simulated using the PYTHIA8 [50] event generator with the injection of signals. The generated signals are transported through the detector using GEANT3 [51]. The is obtained from INEL0 events and applied uniformly across all multiplicity classes. It ranges from 26% at 0 0.3 GeV/ to 60% as increases, and shows no significant dependence on event multiplicity. The of the to decay is estimated to be (46 6)% [25]. The correction for the signal loss (), which compensates for signal losses due to event selection, is derived using the meson as a surrogate because typical Monte Carlo event generators do not generate primary particles. This method, exploiting universal scaling [52], shows that does not depend on particle species [53]. It is quantified at 1.08 for 0 0.3 GeV/ in 50–100% multiplicity class and approaches unity as increases. In the highest multiplicity class, is near unity for the range this analysis measures.
4 Systematic uncertainties
The systematic uncertainties related to the production of are estimated by varying the analysis selection criteria, the fit configuration utilised to extract the raw yield, and the phase space correction treatment. For each systematic variation, a conservative maximal requirement of 1.3 standard deviations is adopted for the Barlow test [54]. Table 1 summarises the estimated uncertainties. The total systematic uncertainty is calculated by adding the different contributing sources in a quadratic sum. For each source, the minimum and maximum uncertainty values are reported in the table. The relative uncertainty of the is 13% [25] and is not included in the total uncertainty.
| Sources | Systematic uncertainty (%) |
| PV selection | negligible |
| Pileup rejection | negligible |
| Tracking | 1.4 |
| Material budget | 0.4–2 |
| ITS–TPC matching | 6 |
| Particle identification | 2–10 |
| (1270) parameters | 1–7 |
| (770)0 parameters | 1–8 |
| Fit range | 0–8 |
| Initial width | 1–8 |
| Phase space correction | 2–8 |
| Total (in quadrature) | 10–18 |
The analysis is repeated with a different PV selection of 7 cm to estimate the systematic uncertainty from the PV selection. The systematic uncertainty is found to be negligible compared to the statistical fluctuation arising from the selection variation. The systematic uncertainty from the pileup rejection is studied by varying the number of track segments used to reconstruct pileup collision vertices from 5 to 3, and the resulting uncertainty is found to be negligible.
The systematic uncertainty on the tracking is determined by varying the requirements used to select reconstructed primary tracks, such as those on , , and the number of crossed rows. According to Ref. [45], this results in an estimated uncertainty of 1.4%. The systematic uncertainty from PID is tested with different requirements on the number of standard deviations ( with respect to the default selections) for the TPC and TOF selection. The resulting uncertainties vary, ranging from 2% to 10% in the different intervals and multiplicity classes.
The systematic uncertainty on the extraction of the yield through fits to the invariant mass distributions is estimated by changing certain configurations in the fit procedure. The contributions from the masses and widths of f2(1270) and (770)0 are evaluated by shifting the masses and widths by three standard deviations with respect to their world-average values, using the uncertainties reported in Ref. [12]. The estimated systematic uncertainties are 1–7% and 1–8%, respectively. In addition, the invariant mass range used in the fit is varied inward or outward by 40 MeV/, and the resulting systematic uncertainty is found to be less than 8%. The systematic uncertainty from the initial width value, which is obtained in the first fit step described in Sec. 3, is estimated by varying the width within the statistical uncertainties, which is about 5 MeV/ on average, in both directions. These variations affect the background distribution determined in the second step, and the estimated systematic uncertainties range from 1% to 8%. The systematic uncertainty from the phase space correction is estimated by varying the kinetic freeze-out temperature in the range of 140 180 MeV. The estimated uncertainties range from 2% to 8%.
The systematic uncertainty from the ITS–TPC matching is estimated by comparing the matching efficiencies in experimental data and MC samples and is 6% [45]. The systematic uncertainty from the material budget is estimated by varying the amount of the budget by 5% in the simulation and is 0.4–2% [45].
The correlations between systematic uncertainties across different multiplicity classes are quantitatively assessed. A stronger correlation is expected when the systematic deviations exhibit the same direction across different multiplicity classes. To test the correlations, the directions of the systematic deviations between a given multiplicity class and the MB class are compared. The systematic uncertainties analysed in this study account for approximately half of the total systematic uncertainties, with the remaining half being uncorrelated.
5 Results
Figure 2 shows the spectra of in pp collisions at 13 TeV measured in the range of 0 13 GeV/ for different multiplicity classes and INEL 0 events, which is defined in Ref [55]. Each spectrum is scaled by a multiplicative factor, denoted in the figure, for better visibility. The lower panel of Fig. 2 shows the ratios of the spectra in different multiplicity classes to the INEL 0 one. The systematic uncertainties of the ratios are calculated by propagating the multiplicity-uncorrelated uncertainties on the individual spectra. For 4 GeV/, a hardening of the spectrum from low- to high-multiplicity events is observed, while the spectral shapes in the different multiplicity classes are found to become consistent among each other for 4 GeV/. Similar trends are observed also for other hadron species [56, 57] and measured in p–Pb collisions [13].
Figure 3 shows the -integrated yield (ddy) and mean transverse momentum () of as a function of charged-particle multiplicity in different collision systems [13, 48]. The d/d and are calculated by integrating and averaging the spectrum, respectively, in different multiplicity classes. The dd of is found to increase linearly with the charged-particle multiplicity, and the increases with multiplicity and then seems to saturate at high multiplicity, as observed for other hadron species [45].
Figure 4 shows the particle yield ratio of to charged pions as a function of charged-particle multiplicity in different collision systems. The decreases with increasing multiplicity across different collision systems. The experimental data are compared with theoretical calculations from the CSM [21]. The CSM calculation with the assumption of 2 for expects to increase with increasing multiplicity because it assumes a specific quark content for the , including strange quarks. On the other hand, the CSM calculation with the assumption of 0 for qualitatively describes the decreasing trend, which is observed in experimental data. However, a quantitative description is not achieved as the model is consistent only with the point in the highest-multiplicity events, and the model provides an underestimation of 39% of the particle yield ratio in the lowest-multiplicity events. The model provides good description of particle yield ratio in low multiplicity events and overestimates the ratio in high multiplicity events for other hadron species [21]. In both the and scenarios, an additional decrease with increasing multiplicity can be expected if the rescattering effect is considered, and in the scenario, the increase could be converted to a decrease. The model prediction for the scenario is closer to the data.
Figure 5 shows the particle yield ratio of to K∗(892)0 as a function of charged-particle multiplicity in different collision systems. The K∗(892)0 decreases with increasing multiplicity in different collision systems. The CSM calculation with the assumption of 0 also predicts the decreasing trend of the particle yield ratio, whereas the calculation with the the scenario does not. However, with the scenario, the model underestimates the particle yield ratio by about 35% throughout the multiplicity range, which is of a similar magnitude to the underestimation observed in the / ratio. Because rescattering affects both the and (892)0 resonances, its effects are expected to partially cancel out in the /(892)0 yield ratio. However, despite their similar lifetimes, they differ in their decay products, where and (892)0 decay into and K, respectively. The rescattering effect for can be stronger than for (892)0 when the scattering cross section is greater than . Similar to ratio, the model prediction for (892)0 with the scenario is closer to the data.
Figure 6 presents the -integrated particle yield double ratio of to (left panel) and (892)0-(right panel) as a function of charged-particle multiplicity in pp collisions at TeV. The double ratio is defined as the ratio of the to hadron yield ratio in a given multiplicity class to the corresponding value measured in INEL 0 events. This approach facilitates a direct comparison of the multiplicity dependence while mitigating correlated systematic uncertainties and cancelling uncertainties related to branching ratios. For the double ratios, the CSM prediction with the scenario is closer to the data.
Figure 7 shows the -differential particle yield ratio of to (left panel) and (892)0 (right panel) as a function of charged-particle multiplicity in pp collisions at TeV, along with their double ratios relative to INEL 0 events in the lower panels. For better visibility, the single ratios are scaled by the factors indicated in the legend. The systematic uncertainties of the double ratios in the lower panels account only for uncorrelated uncertainties. The ratio of to increases with across all measured multiplicity classes. The double ratio at low multiplicity is higher than that at high multiplicity, particularly at low , and approaches unity at higher , suggesting that is relatively suppressed only at low . On the other hand, the ratio of to (892)0 is nearly constant with , likely due to the similar mass for both particles. The double ratio of to (892)0 shows a mild decrease with increasing multiplicity but exhibits no significant dependence, indicating that the suppression mechanism does not vary with transverse momentum.
6 Conclusions
The multiplicity and dependence of production in pp collisions at = 13 TeV is presented. The is reconstructed via the decay channel at midrapidity ( 0.5) in the transverse momentum region of 13 GeV/. A hardening of the spectra and a consequent increase of the mean are observed with increasing multiplicity.
The -integrated particle yield ratio of to decreases with increasing multiplicity. The CSM calculation shows qualitatively such a trend for the scenario but yields to an unexpected underestimation of the ratio in low-multiplicity events. The -differential particle yield ratio shows that is relatively suppressed only at low . These phenomena are also observed for short-lived resonances produced in heavy-ion collisions, where it is attributed to be the result of hadronic rescatterings.
The -integrated particle yield ratio of to (892)0 decreases with increasing multiplicity, as expected under the scenario. In addition, the CSM calculation underestimates the ratio across the measured multiplicity range. However, yields of two particles can be modified differently by hadronic interactions, which complicates the physical interpretation. This ambiguity could be reduced by measuring the particle yield ratio in large collision systems, where the strangeness enhancement is no longer present and only hadronic interaction effects remain.
The comparison between the data and model calculations suggests that the is less likely to contain a strange or antistrange quark, and therefore is unlikely to have a tetraquark structure. Given the large uncertainty in the branching ratio of and the lack of differential model calculations for hadronic interactions on the , further developments in these aspects would greatly improve our understanding of the .
Acknowledgements
The ALICE Collaboration would like to thank all its engineers and technicians for their invaluable contributions to the construction of the experiment and the CERN accelerator teams for the outstanding performance of the LHC complex. The ALICE Collaboration gratefully acknowledges the resources and support provided by all Grid centres and the Worldwide LHC Computing Grid (WLCG) collaboration. The ALICE Collaboration acknowledges the following funding agencies for their support in building and running the ALICE detector: A. I. Alikhanyan National Science Laboratory (Yerevan Physics Institute) Foundation (ANSL), State Committee of Science and World Federation of Scientists (WFS), Armenia; Austrian Academy of Sciences, Austrian Science Fund (FWF): [M 2467-N36] and Nationalstiftung für Forschung, Technologie und Entwicklung, Austria; Ministry of Communications and High Technologies, National Nuclear Research Center, Azerbaijan; Rede Nacional de Física de Altas Energias (Renafae), Financiadora de Estudos e Projetos (Finep), Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) and The Sao Paulo Research Foundation (FAPESP), Brazil; Bulgarian Ministry of Education and Science, within the National Roadmap for Research Infrastructures 2020-2027 (object CERN), Bulgaria; Ministry of Education of China (MOEC) , Ministry of Science & Technology of China (MSTC) and National Natural Science Foundation of China (NSFC), China; Ministry of Science and Education and Croatian Science Foundation, Croatia; Centro de Aplicaciones Tecnológicas y Desarrollo Nuclear (CEADEN), Cubaenergía, Cuba; Ministry of Education, Youth and Sports of the Czech Republic, Czech Republic; The Danish Council for Independent Research | Natural Sciences, the VILLUM FONDEN and Danish National Research Foundation (DNRF), Denmark; Helsinki Institute of Physics (HIP), Finland; Commissariat à l’Energie Atomique (CEA) and Institut National de Physique Nucléaire et de Physique des Particules (IN2P3) and Centre National de la Recherche Scientifique (CNRS), France; Bundesministerium für Bildung und Forschung (BMBF) and GSI Helmholtzzentrum für Schwerionenforschung GmbH, Germany; General Secretariat for Research and Technology, Ministry of Education, Research and Religions, Greece; National Research, Development and Innovation Office, Hungary; Department of Atomic Energy Government of India (DAE), Department of Science and Technology, Government of India (DST), University Grants Commission, Government of India (UGC) and Council of Scientific and Industrial Research (CSIR), India; National Research and Innovation Agency - BRIN, Indonesia; Istituto Nazionale di Fisica Nucleare (INFN), Italy; Japanese Ministry of Education, Culture, Sports, Science and Technology (MEXT) and Japan Society for the Promotion of Science (JSPS) KAKENHI, Japan; Consejo Nacional de Ciencia (CONACYT) y Tecnología, through Fondo de Cooperación Internacional en Ciencia y Tecnología (FONCICYT) and Dirección General de Asuntos del Personal Academico (DGAPA), Mexico; Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO), Netherlands; The Research Council of Norway, Norway; Pontificia Universidad Católica del Perú, Peru; Ministry of Science and Higher Education, National Science Centre and WUT ID-UB, Poland; Korea Institute of Science and Technology Information and National Research Foundation of Korea (NRF), Republic of Korea; Ministry of Education and Scientific Research, Institute of Atomic Physics, Ministry of Research and Innovation and Institute of Atomic Physics and Universitatea Nationala de Stiinta si Tehnologie Politehnica Bucuresti, Romania; Ministerstvo skolstva, vyskumu, vyvoja a mladeze SR, Slovakia; National Research Foundation of South Africa, South Africa; Swedish Research Council (VR) and Knut & Alice Wallenberg Foundation (KAW), Sweden; European Organization for Nuclear Research, Switzerland; Suranaree University of Technology (SUT), National Science and Technology Development Agency (NSTDA) and National Science, Research and Innovation Fund (NSRF via PMU-B B05F650021), Thailand; Turkish Energy, Nuclear and Mineral Research Agency (TENMAK), Turkey; National Academy of Sciences of Ukraine, Ukraine; Science and Technology Facilities Council (STFC), United Kingdom; National Science Foundation of the United States of America (NSF) and United States Department of Energy, Office of Nuclear Physics (DOE NP), United States of America. In addition, individual groups or members have received support from: Czech Science Foundation (grant no. 23-07499S), Czech Republic; FORTE project, reg. no. CZ.02.01.01/00/22_008/0004632, Czech Republic, co-funded by the European Union, Czech Republic; European Research Council (grant no. 950692), European Union; Deutsche Forschungs Gemeinschaft (DFG, German Research Foundation) “Neutrinos and Dark Matter in Astro- and Particle Physics” (grant no. SFB 1258), Germany; ICSC - National Research Center for High Performance Computing, Big Data and Quantum Computing and FAIR - Future Artificial Intelligence Research, funded by the NextGenerationEU program (Italy).
Appendix A The ALICE Collaboration
I.J. Abualrob 114, S. Acharya 50, G. Aglieri Rinella 32, L. Aglietta 24, M. Agnello 29, N. Agrawal 25, Z. Ahammed 133, S. Ahmad 15, I. Ahuja 36, ZUL. Akbar81, A. Akindinov 139, V. Akishina38, M. Al-Turany 96, D. Aleksandrov 139, B. Alessandro 56, H.M. Alfanda 6, R. Alfaro Molina 67, B. Ali 15, A. Alici 25, A. Alkin 103, J. Alme 20, G. Alocco 24, T. Alt 64, A.R. Altamura 50, I. Altsybeev 94, C. Andrei 45, N. Andreou 113, A. Andronic 124, E. Andronov 139, V. Anguelov 93, F. Antinori 54, P. Antonioli 51, N. Apadula 73, H. Appelshäuser 64, C. Arata 72, S. Arcelli 25, R. Arnaldi 56, J.G.M.C.A. Arneiro 109, I.C. Arsene 19, M. Arslandok 136, A. Augustinus 32, R. Averbeck 96, M.D. Azmi 15, H. Baba122, A.R.J. Babu135, A. Badalà 53, J. Bae 103, Y. Bae 103, Y.W. Baek 40, X. Bai 118, R. Bailhache 64, Y. Bailung 48, R. Bala 90, A. Baldisseri 128, B. Balis 2, S. Bangalia116, Z. Banoo 90, V. Barbasova 36, F. Barile 31, L. Barioglio 56, M. Barlou 24,77, B. Barman 41, G.G. Barnaföldi 46, L.S. Barnby 113, E. Barreau 102, V. Barret 125, L. Barreto 109, K. Barth 32, E. Bartsch 64, N. Bastid 125, G. Batigne 102, D. Battistini 94, B. Batyunya 140, D. Bauri47, J.L. Bazo Alba 100, I.G. Bearden 82, P. Becht 96, D. Behera 48, S. Behera 47, I. Belikov 127, V.D. Bella 127, F. Bellini 25, R. Bellwied 114, L.G.E. Beltran 108, Y.A.V. Beltran 44, G. Bencedi 46, A. Bensaoula114, S. Beole 24, Y. Berdnikov 139, A. Berdnikova 93, L. Bergmann 73,93, L. Bernardinis 23, L. Betev 32, P.P. Bhaduri 133, T. Bhalla89, A. Bhasin 90, B. Bhattacharjee 41, S. Bhattarai116, L. Bianchi 24, J. Bielčík 34, J. Bielčíková 85, A. Bilandzic 94, A. Binoy 116, G. Biro 46, S. Biswas 4, D. Blau 139, M.B. Blidaru 96, N. Bluhme38, C. Blume 64, F. Bock 86, T. Bodova 20, J. Bok 16, L. Boldizsár 46, M. Bombara 36, P.M. Bond 32, G. Bonomi 132,55, H. Borel 128, A. Borissov 139, A.G. Borquez Carcamo 93, E. Botta 24, Y.E.M. Bouziani 64, D.C. Brandibur 63, L. Bratrud 64, P. Braun-Munzinger 96, M. Bregant 109, M. Broz 34, G.E. Bruno 95,31, V.D. Buchakchiev 35, M.D. Buckland 84, H. Buesching 64, S. Bufalino 29, P. Buhler 101, N. Burmasov 140, Z. Buthelezi 68,121, A. Bylinkin 20, C. Carr 99, J.C. Cabanillas Noris 108, M.F.T. Cabrera 114, H. Caines 136, A. Caliva 28, E. Calvo Villar 100, J.M.M. Camacho 108, P. Camerini 23, M.T. Camerlingo 50, F.D.M. Canedo 109, S. Cannito 23, S.L. Cantway 136, M. Carabas 112, F. Carnesecchi 32, L.A.D. Carvalho 109, J. Castillo Castellanos 128, M. Castoldi 32, F. Catalano 32, S. Cattaruzzi 23, R. Cerri 24, I. Chakaberia 73, P. Chakraborty 134, J.W.O. ChanI,114, S. Chandra 133, S. Chapeland 32, M. Chartier 117, S. Chattopadhay133, M. Chen 39, T. Cheng 6, C. Cheshkov 126, D. Chiappara 27, V. Chibante Barroso 32, D.D. Chinellato 101, F. Chinu 24, E.S. Chizzali II,94, J. Cho 58, S. Cho 58, P. Chochula 32, Z.A. Chochulska III,134, D. Choudhury41, P. Christakoglou 83, C.H. Christensen 82, P. Christiansen 74, T. Chujo 123, M. Ciacco 29, C. Cicalo 52, G. Cimador 24, F. Cindolo 51, M.R. Ciupek96, G. ClaiIV,51, F. Colamaria 50, J.S. Colburn99, D. Colella 31, A. Colelli31, M. Colocci 25, M. Concas 32, G. Conesa Balbastre 72, Z. Conesa del Valle 129, G. Contin 23, J.G. Contreras 34, M.L. Coquet 102, P. Cortese 131,56, M.R. Cosentino 111, F. Costa 32, S. Costanza 21, P. Crochet 125, M.M. Czarnynoga134, A. Dainese 54, G. Dange38, M.C. Danisch 93, A. Danu 63, P. Das 32, S. Das 4, A.R. Dash 124, S. Dash 47, A. De Caro 28, G. de Cataldo 50, J. de Cuveland 38, A. De Falco 22, D. De Gruttola 28, N. De Marco 56, C. De Martin 23, S. De Pasquale 28, R. Deb 132, R. Del Grande 94, L. Dello Stritto 32, G.G.A. de Souza V,109, P. Dhankher 18, D. Di Bari 31, M. Di Costanzo 29, A. Di Mauro 32, B. Di Ruzza 130, B. Diab 32, Y. Ding 6, J. Ditzel 64, R. Divià 32, Ø. Djuvsland20, A. Dobrin 63, B. Dönigus 64, L. Döpper 42, J.M. Dubinski 134, A. Dubla 96, P. Dupieux 125, N. Dzalaiova13, T.M. Eder 124, R.J. Ehlers 73, F. Eisenhut 64, R. Ejima 91, D. Elia 50, B. Erazmus 102, F. Ercolessi 25, B. Espagnon 129, G. Eulisse 32, D. Evans 99, L. Fabbietti 94, M. Faggin 32, J. Faivre 72, F. Fan 6, W. Fan 73, T. Fang6, A. Fantoni 49, M. Fasel 86, G. Feofilov 139, A. Fernández Téllez 44, L. Ferrandi 109, M.B. Ferrer 32, A. Ferrero 128, C. Ferrero VI,56, A. Ferretti 24, V.J.G. Feuillard 93, D. Finogeev 140, F.M. Fionda 52, A.N. Flores 107, S. Foertsch 68, I. Fokin 93, S. Fokin 139, U. Follo VI,56, R. Forynski 113, E. Fragiacomo 57, E. Frajna 46, H. Fribert 94, U. Fuchs 32, N. Funicello 28, C. Furget 72, A. Furs 140, T. Fusayasu 97, J.J. Gaardhøje 82, M. Gagliardi 24, A.M. Gago 100, T. Gahlaut47, C.D. Galvan 108, S. Gami 79, D.R. Gangadharan 114, P. Ganoti 77, C. Garabatos 96, J.M. Garcia 44, T. García Chávez 44, E. Garcia-Solis 9, S. Garetti 129, C. Gargiulo 32, P. Gasik 96, H.M. Gaur38, A. Gautam 116, M.B. Gay Ducati 66, M. Germain 102, R.A. Gernhaeuser 94, C. Ghosh133, M. Giacalone 51, G. Gioachin 29, S.K. Giri 133, P. Giubellino 96,56, P. Giubilato 27, P. Glässel 93, E. Glimos 120, V. Gonzalez 135, M. Gorgon 2, K. Goswami 48, S. Gotovac 33, V. Grabski 67, L.K. Graczykowski 134, E. Grecka 85, A. Grelli 59, C. Grigoras 32, V. Grigoriev 139, S. Grigoryan 140,1, O.S. Groettvik 32, F. Grosa 32, J.F. Grosse-Oetringhaus 32, R. Grosso 96, D. Grund 34, N.A. Grunwald 93, R. Guernane 72, M. Guilbaud 102, K. Gulbrandsen 82, J.K. Gumprecht 101, T. Gündem 64, T. Gunji 122, J. Guo10, W. Guo 6, A. Gupta 90, R. Gupta 90, R. Gupta 48, K. Gwizdziel 134, L. Gyulai 46, C. Hadjidakis 129, F.U. Haider 90, S. Haidlova 34, M. Haldar4, H. Hamagaki 75, Y. Han 138, B.G. Hanley 135, R. Hannigan 107, J. Hansen 74, J.W. Harris 136, A. Harton 9, M.V. Hartung 64, H. Hassan 115, D. Hatzifotiadou 51, P. Hauer 42, L.B. Havener 136, E. Hellbär 32, H. Helstrup 37, M. Hemmer 64, T. Herman 34, S.G. Hernandez114, G. Herrera Corral 8, K.F. Hetland 37, B. Heybeck 64, H. Hillemanns 32, B. Hippolyte 127, I.P.M. Hobus 83, F.W. Hoffmann 70, B. Hofman 59, M. Horst 94, A. Horzyk 2, Y. Hou 96,11,6, P. Hristov 32, P. Huhn64, L.M. Huhta 115, T.J. Humanic 87, V. Humlova 34, A. Hutson 114, D. Hutter 38, M.C. Hwang 18, R. Ilkaev139, M. Inaba 123, M. Ippolitov 139, A. Isakov 83, T. Isidori 116, M.S. Islam 47, M. Ivanov13, M. Ivanov 96, K.E. Iversen 74, J.G.Kim 138, M. Jablonski 2, B. Jacak 18,73, N. Jacazio 25, P.M. Jacobs 73, S. Jadlovska105, J. Jadlovsky105, S. Jaelani 81, C. Jahnke 110, M.J. Jakubowska 134, E.P. Jamro 2, D.M. Janik 34, M.A. Janik 134, S. Ji 16, S. Jia 82, T. Jiang 10, A.A.P. Jimenez 65, S. JinI,10, F. Jonas 73, D.M. Jones 117, J.M. Jowett 32,96, J. Jung 64, M. Jung 64, A. Junique 32, A. Jusko 99, J. Kaewjai104, P. Kalinak 60, A. Kalweit 32, A. Karasu Uysal 137, N. Karatzenis99, O. Karavichev 139, T. Karavicheva 139, M.J. Karwowska 134, U. Kebschull 70, M. Keil 32, B. Ketzer 42, J. Keul 64, S.S. Khade 48, A.M. Khan 118, A. Khanzadeev 139, Y. Kharlov 139, A. Khatun 116, A. Khuntia 51, Z. Khuranova 64, B. Kileng 37, B. Kim 103, C. Kim 16, D.J. Kim 115, D. Kim 103, E.J. Kim 69, G. Kim 58, H. Kim 58, J. Kim 138, J. Kim 58, J. Kim 32, M. Kim 18, S. Kim 17, T. Kim 138, K. Kimura 91, S. Kirsch 64, I. Kisel 38, S. Kiselev 139, A. Kisiel 134, J.L. Klay 5, J. Klein 32, S. Klein 73, C. Klein-Bösing 124, M. Kleiner 64, A. Kluge 32, C. Kobdaj 104, R. Kohara 122, T. Kollegger96, A. Kondratyev 140, N. Kondratyeva 139, J. Konig 64, P.J. Konopka 32, G. Kornakov 134, M. Korwieser 94, S.D. Koryciak 2, C. Koster 83, A. Kotliarov 85, N. Kovacic 88, V. Kovalenko 139, M. Kowalski 106, V. Kozhuharov 35, G. Kozlov 38, I. Králik 60, A. Kravčáková 36, L. Krcal 32, M. Krivda 99,60, F. Krizek 85, K. Krizkova Gajdosova 34, C. Krug 66, E. Kryshen 139, V. Kučera 58, C. Kuhn 127, T. Kumaoka123, D. Kumar 133, L. Kumar 89, N. Kumar89, S. Kumar 50, S. Kundu 32, M. Kuo123, P. Kurashvili 78, A.B. Kurepin 139, S. Kurita 91, A. Kuryakin 139, S. Kushpil 85, M. Kutyla134, A. Kuznetsov 140, M.J. Kweon 58, Y. Kwon 138, S.L. La Pointe 38, P. La Rocca 26, A. Lakrathok104, M. Lamanna 32, S. Lambert102, A.R. Landou 72, R. Langoy 119, P. Larionov 32, E. Laudi 32, L. Lautner 94, R.A.N. Laveaga 108, R. Lavicka 101, R. Lea 132,55, H. Lee 103, I. Legrand 45, G. Legras 124, A.M. Lejeune 34, T.M. Lelek 2, R.C. Lemmon I,84, I. León Monzón 108, M.M. Lesch 94, P. Lévai 46, M. Li6, P. Li10, X. Li10, B.E. Liang-Gilman 18, J. Lien 119, R. Lietava 99, I. Likmeta 114, B. Lim 56, H. Lim 16, S.H. Lim 16, S. Lin10, V. Lindenstruth 38, C. Lippmann 96, D. Liskova 105, D.H. Liu 6, J. Liu 117, G.S.S. Liveraro 110, I.M. Lofnes 20, C. Loizides 86, S. Lokos 106, J. Lömker 59, X. Lopez 125, E. López Torres 7, C. Lotteau 126, P. Lu 96,118, W. Lu 6, Z. Lu 10, F.V. Lugo 67, J. Luo39, G. Luparello 57, M.A.T. Johnson 44, Y.G. Ma 39, M. Mager 32, A. Maire 127, E.M. Majerz 2, M.V. Makariev 35, G. Malfattore 51, N.M. Malik 90, N. Malik 15, S.K. Malik 90, D. Mallick 129, N. Mallick 115, G. Mandaglio 30,53, S.K. Mandal 78, A. Manea 63, V. Manko 139, A.K. Manna48, F. Manso 125, G. Mantzaridis 94, V. Manzari 50, Y. Mao 6, R.W. Marcjan 2, G.V. Margagliotti 23, A. Margotti 51, A. Marín 96, C. Markert 107, P. Martinengo 32, M.I. Martínez 44, G. Martínez García 102, M.P.P. Martins 32,109, S. Masciocchi 96, M. Masera 24, A. Masoni 52, L. Massacrier 129, O. Massen 59, A. Mastroserio 130,50, L. Mattei 24,125, S. Mattiazzo 27, A. Matyja 106, F. Mazzaschi 32, M. Mazzilli 31,114, Y. Melikyan 43, M. Melo 109, A. Menchaca-Rocha 67, J.E.M. Mendez 65, E. Meninno 101, A.S. Menon 114, M.W. Menzel32,93, M. Meres 13, L. Micheletti 56, D. Mihai112, D.L. Mihaylov 94, A.U. Mikalsen 20, K. Mikhaylov 140,139, L. Millot 72, N. Minafra 116, D. Miśkowiec 96, A. Modak 57,132, B. Mohanty 79, M. Mohisin Khan VII,15, M.A. Molander 43, M.M. Mondal 79, S. Monira 134, D.A. Moreira De Godoy 124, A. Morsch 32, T. Mrnjavac 32, S. Mrozinski 64, V. Muccifora 49, S. Muhuri 133, A. Mulliri 22, M.G. Munhoz 109, R.H. Munzer 64, H. Murakami 122, L. Musa 32, J. Musinsky 60, J.W. Myrcha 134, N.B.Sundstrom 59, B. Naik 121, A.I. Nambrath 18, B.K. Nandi 47, R. Nania 51, E. Nappi 50, A.F. Nassirpour 17, V. Nastase112, A. Nath 93, N.F. Nathanson 82, C. Nattrass 120, K. Naumov18, A. Neagu19, L. Nellen 65, R. Nepeivoda 74, S. Nese 19, N. Nicassio 31, B.S. Nielsen 82, E.G. Nielsen 82, S. Nikolaev 139, V. Nikulin 139, F. Noferini 51, S. Noh 12, P. Nomokonov 140, J. Norman 117, N. Novitzky 86, J. Nystrand 20, M.R. Ockleton117, M. Ogino 75, S. Oh 17, A. Ohlson 74, M. Oida 91, V.A. Okorokov 139, J. Oleniacz 134, C. Oppedisano 56, A. Ortiz Velasquez 65, H. Osanai75, J. Otwinowski 106, M. Oya91, K. Oyama 75, S. Padhan 47, D. Pagano 132,55, G. Paić 65, S. Paisano-Guzmán 44, A. Palasciano 50, I. Panasenko 74, S. Panebianco 128, P. Panigrahi 47, C. Pantouvakis 27, H. Park 123, J. Park 123, S. Park 103, T.Y. Park138, J.E. Parkkila 134, P.B. Pati 82, Y. Patley 47, R.N. Patra50, P. Paudel116, B. Paul 133, H. Pei 6, T. Peitzmann 59, X. Peng 11, M. Pennisi 24, S. Perciballi 24, D. Peresunko 139, G.M. Perez 7, Y. Pestov139, M. Petrovici 45, S. Piano 57, M. Pikna 13, P. Pillot 102, O. Pinazza 51,32, L. Pinsky114, C. Pinto 32, S. Pisano 49, M. Płoskoń 73, M. Planinic 88, D.K. Plociennik 2, M.G. Poghosyan 86, B. Polichtchouk 139, S. Politano 32,24, N. Poljak 88, A. Pop 45, S. Porteboeuf-Houssais 125, I.Y. Pozos 44, K.K. Pradhan 48, S.K. Prasad 4, S. Prasad 48, R. Preghenella 51, F. Prino 56, C.A. Pruneau 135, I. Pshenichnov 139, M. Puccio 32, S. Pucillo 28,24, L. Quaglia 24, A.M.K. Radhakrishnan 48, S. Ragoni 14, A. Rai 136, A. Rakotozafindrabe 128, N. Ramasubramanian126, L. Ramello 131,56, C.O. Ramírez-Álvarez 44, M. Rasa 26, S.S. Räsänen 43, R. Rath 96, M.P. Rauch 20, I. Ravasenga 32, K.F. Read 86,120, C. Reckziegel 111, A.R. Redelbach 38, K. Redlich VIII,78, C.A. Reetz 96, H.D. Regules-Medel 44, A. Rehman 20, F. Reidt 32, H.A. Reme-Ness 37, K. Reygers 93, R. Ricci 28, M. Richter 20, A.A. Riedel 94, W. Riegler 32, A.G. Riffero 24, M. Rignanese 27, C. Ripoli 28, C. Ristea 63, M.V. Rodriguez 32, M. Rodríguez Cahuantzi 44, K. Røed 19, R. Rogalev 139, E. Rogochaya 140, D. Rohr 32, D. Röhrich 20, S. Rojas Torres 34, P.S. Rokita 134, G. Romanenko 25, F. Ronchetti 32, D. Rosales Herrera 44, E.D. Rosas65, K. Roslon 134, A. Rossi 54, A. Roy 48, S. Roy 47, N. Rubini 51, J.A. Rudolph83, D. Ruggiano 134, R. Rui 23, P.G. Russek 2, R. Russo 83, A. Rustamov 80, Y. Ryabov 139, A. Rybicki 106, L.C.V. Ryder 116, G. Ryu 71, J. Ryu 16, W. Rzesa 134, B. Sabiu 51, R. Sadek 73, S. Sadhu 42, S. Sadovsky 139, S. Saha 79, B. Sahoo 48, R. Sahoo 48, D. Sahu 48, P.K. Sahu 61, J. Saini 133, K. Sajdakova36, S. Sakai 123, S. Sambyal 90, D. Samitz 101, I. Sanna 32,94, T.B. Saramela109, D. Sarkar 82, P. Sarma 41, V. Sarritzu 22, V.M. Sarti 94, M.H.P. Sas 32, S. Sawan 79, E. Scapparone 51, J. Schambach 86, H.S. Scheid 32, C. Schiaua 45, R. Schicker 93, F. Schlepper 32,93, A. Schmah96, C. Schmidt 96, M.O. Schmidt 32, M. Schmidt92, N.V. Schmidt 86, A.R. Schmier 120, J. Schoengarth 64, R. Schotter 101, A. Schröter 38, J. Schukraft 32, K. Schweda 96, G. Scioli 25, E. Scomparin 56, J.E. Seger 14, Y. Sekiguchi122, D. Sekihata 122, M. Selina 83, I. Selyuzhenkov 96, S. Senyukov 127, J.J. Seo 93, D. Serebryakov 139, L. Serkin IX,65, L. Šerkšnytė 94, A. Sevcenco 63, T.J. Shaba 68, A. Shabetai 102, R. Shahoyan 32, B. Sharma 90, D. Sharma 47, H. Sharma 54, M. Sharma 90, S. Sharma 90, T. Sharma 41, U. Sharma 90, A. Shatat 129, O. Sheibani135, K. Shigaki 91, M. Shimomura 76, S. Shirinkin 139, Q. Shou 39, Y. Sibiriak 139, S. Siddhanta 52, T. Siemiarczuk 78, T.F. Silva 109, W.D. Silva I,109, D. Silvermyr 74, T. Simantathammakul 104, R. Simeonov 35, B. Singh90, B. Singh 94, K. Singh 48, R. Singh 79, R. Singh 54,96, S. Singh 15, V.K. Singh 133, V. Singhal 133, T. Sinha 98, B. Sitar 13, M. Sitta 131,56, T.B. Skaali 19, G. Skorodumovs 93, N. Smirnov 136, R.J.M. Snellings 59, E.H. Solheim 19, C. Sonnabend 32,96, J.M. Sonneveld 83, F. Soramel 27, A.B. Soto-Hernandez 87, R. Spijkers 83, I. Sputowska 106, J. Staa 74, J. Stachel 93, I. Stan 63, T. Stellhorn 124, S.F. Stiefelmaier 93, D. Stocco 102, I. Storehaug 19, N.J. Strangmann 64, P. Stratmann 124, S. Strazzi 25, A. Sturniolo 30,53, A.A.P. Suaide 109, C. Suire 129, A. Suiu 32,112, M. Sukhanov 140, M. Suljic 32, R. Sultanov 139, V. Sumberia 90, S. Sumowidagdo 81, L.H. Tabares 7, S.F. Taghavi 94, J. Takahashi 110, G.J. Tambave 79, Z. Tang 118, J. Tanwar 89, J.D. Tapia Takaki 116, N. Tapus 112, L.A. Tarasovicova 36, M.G. Tarzila 45, A. Tauro 32, A. Tavira García 129, G. Tejeda Muñoz 44, L. Terlizzi 24, C. Terrevoli 50, D. Thakur 24, S. Thakur 4, M. Thogersen 19, D. Thomas 107, N. Tiltmann 32,124, A.R. Timmins 114, A. Toia 64, R. Tokumoto91, S. Tomassini 25, K. Tomohiro91, N. Topilskaya 139, M. Toppi 49, V.V. Torres 102, A. Trifiró 30,53, T. Triloki 95, A.S. Triolo 32,53, S. Tripathy 32, T. Tripathy 125, S. Trogolo 24, V. Trubnikov 3, W.H. Trzaska 115, T.P. Trzcinski 134, C. Tsolanta19, R. Tu39, A. Tumkin 139, R. Turrisi 54, T.S. Tveter 19, K. Ullaland 20, B. Ulukutlu 94, S. Upadhyaya 106, A. Uras 126, M. Urioni 23, G.L. Usai 22, M. Vaid90, M. Vala 36, N. Valle 55, L.V.R. van Doremalen59, M. van Leeuwen 83, C.A. van Veen 93, R.J.G. van Weelden 83, D. Varga 46, Z. Varga 136, P. Vargas Torres65, M. Vasileiou 77, A. Vasiliev I,139, O. Vázquez Doce 49, O. Vazquez Rueda 114, V. Vechernin 139, P. Veen 128, E. Vercellin 24, R. Verma 47, R. Vértesi 46, M. Verweij 59, L. Vickovic33, Z. Vilakazi121, O. Villalobos Baillie 99, A. Villani 23, A. Vinogradov 139, T. Virgili 28, M.M.O. Virta 115, A. Vodopyanov 140, B. Volkel 32, M.A. Völkl 99, S.A. Voloshin 135, G. Volpe 31, B. von Haller 32, I. Vorobyev 32, N. Vozniuk 140, J. Vrláková 36, J. Wan39, C. Wang 39, D. Wang 39, Y. Wang 39, Y. Wang 6, Z. Wang 39, A. Wegrzynek 32, F. Weiglhofer 32,38, S.C. Wenzel 32, J.P. Wessels 124, P.K. Wiacek 2, J. Wiechula 64, J. Wikne 19, G. Wilk 78, J. Wilkinson 96, G.A. Willems 124, B. Windelband 93, M. Winn 128, J. Witte 96, M. Wojnar 2, J.R. Wright 107, C.-T. Wu 6,27, W. Wu39, Y. Wu 118, K. Xiong39, Z. Xiong118, L. Xu 126,6, R. Xu 6, A. Yadav 42, A.K. Yadav 133, Y. Yamaguchi 91, S. Yang 58, S. Yang 20, S. Yano 91, E.R. Yeats18, J. Yi 6, R. Yin39, Z. Yin 6, I.-K. Yoo 16, J.H. Yoon 58, H. Yu 12, S. Yuan20, A. Yuncu 93, V. Zaccolo 23, C. Zampolli 32, F. Zanone 93, N. Zardoshti 32, P. Závada 62, B. Zhang 93, C. Zhang 128, L. Zhang 39, M. Zhang 125,6, M. Zhang 27,6, S. Zhang 39, X. Zhang 6, Y. Zhang118, Y. Zhang 118, Z. Zhang 6, M. Zhao 10, V. Zherebchevskii 139, Y. Zhi10, D. Zhou 6, Y. Zhou 82, J. Zhu 39, S. Zhu96,118, Y. Zhu6, A. Zingaretti I,54, S.C. Zugravel 56, N. Zurlo 132,55
Affiliation Notes
I Deceased
II Also at: Max-Planck-Institut fur Physik, Munich, Germany
III Also at: Czech Technical University in Prague (CZ)
IV Also at: Italian National Agency for New Technologies, Energy and Sustainable Economic Development (ENEA), Bologna, Italy
V Also at: Instituto de Fisica da Universidade de Sao Paulo
VI Also at: Dipartimento DET del Politecnico di Torino, Turin, Italy
VII Also at: Department of Applied Physics, Aligarh Muslim University, Aligarh, India
VIII Also at: Institute of Theoretical Physics, University of Wroclaw, Poland
IX Also at: Facultad de Ciencias, Universidad Nacional Autónoma de México, Mexico City, Mexico
Collaboration Institutes
1 A.I. Alikhanyan National Science Laboratory (Yerevan Physics Institute) Foundation, Yerevan, Armenia
2 AGH University of Krakow, Cracow, Poland
3 Bogolyubov Institute for Theoretical Physics, National Academy of Sciences of Ukraine, Kyiv, Ukraine
4 Bose Institute, Department of Physics and Centre for Astroparticle Physics and Space Science (CAPSS), Kolkata, India
5 California Polytechnic State University, San Luis Obispo, California, United States
6 Central China Normal University, Wuhan, China
7 Centro de Aplicaciones Tecnológicas y Desarrollo Nuclear (CEADEN), Havana, Cuba
8 Centro de Investigación y de Estudios Avanzados (CINVESTAV), Mexico City and Mérida, Mexico
9 Chicago State University, Chicago, Illinois, United States
10 China Nuclear Data Center, China Institute of Atomic Energy, Beijing, China
11 China University of Geosciences, Wuhan, China
12 Chungbuk National University, Cheongju, Republic of Korea
13 Comenius University Bratislava, Faculty of Mathematics, Physics and Informatics, Bratislava, Slovak Republic
14 Creighton University, Omaha, Nebraska, United States
15 Department of Physics, Aligarh Muslim University, Aligarh, India
16 Department of Physics, Pusan National University, Pusan, Republic of Korea
17 Department of Physics, Sejong University, Seoul, Republic of Korea
18 Department of Physics, University of California, Berkeley, California, United States
19 Department of Physics, University of Oslo, Oslo, Norway
20 Department of Physics and Technology, University of Bergen, Bergen, Norway
21 Dipartimento di Fisica, Università di Pavia, Pavia, Italy
22 Dipartimento di Fisica dell’Università and Sezione INFN, Cagliari, Italy
23 Dipartimento di Fisica dell’Università and Sezione INFN, Trieste, Italy
24 Dipartimento di Fisica dell’Università and Sezione INFN, Turin, Italy
25 Dipartimento di Fisica e Astronomia dell’Università and Sezione INFN, Bologna, Italy
26 Dipartimento di Fisica e Astronomia dell’Università and Sezione INFN, Catania, Italy
27 Dipartimento di Fisica e Astronomia dell’Università and Sezione INFN, Padova, Italy
28 Dipartimento di Fisica ‘E.R. Caianiello’ dell’Università and Gruppo Collegato INFN, Salerno, Italy
29 Dipartimento DISAT del Politecnico and Sezione INFN, Turin, Italy
30 Dipartimento di Scienze MIFT, Università di Messina, Messina, Italy
31 Dipartimento Interateneo di Fisica ‘M. Merlin’ and Sezione INFN, Bari, Italy
32 European Organization for Nuclear Research (CERN), Geneva, Switzerland
33 Faculty of Electrical Engineering, Mechanical Engineering and Naval Architecture, University of Split, Split, Croatia
34 Faculty of Nuclear Sciences and Physical Engineering, Czech Technical University in Prague, Prague, Czech Republic
35 Faculty of Physics, Sofia University, Sofia, Bulgaria
36 Faculty of Science, P.J. Šafárik University, Košice, Slovak Republic
37 Faculty of Technology, Environmental and Social Sciences, Bergen, Norway
38 Frankfurt Institute for Advanced Studies, Johann Wolfgang Goethe-Universität Frankfurt, Frankfurt, Germany
39 Fudan University, Shanghai, China
40 Gangneung-Wonju National University, Gangneung, Republic of Korea
41 Gauhati University, Department of Physics, Guwahati, India
42 Helmholtz-Institut für Strahlen- und Kernphysik, Rheinische Friedrich-Wilhelms-Universität Bonn, Bonn, Germany
43 Helsinki Institute of Physics (HIP), Helsinki, Finland
44 High Energy Physics Group, Universidad Autónoma de Puebla, Puebla, Mexico
45 Horia Hulubei National Institute of Physics and Nuclear Engineering, Bucharest, Romania
46 HUN-REN Wigner Research Centre for Physics, Budapest, Hungary
47 Indian Institute of Technology Bombay (IIT), Mumbai, India
48 Indian Institute of Technology Indore, Indore, India
49 INFN, Laboratori Nazionali di Frascati, Frascati, Italy
50 INFN, Sezione di Bari, Bari, Italy
51 INFN, Sezione di Bologna, Bologna, Italy
52 INFN, Sezione di Cagliari, Cagliari, Italy
53 INFN, Sezione di Catania, Catania, Italy
54 INFN, Sezione di Padova, Padova, Italy
55 INFN, Sezione di Pavia, Pavia, Italy
56 INFN, Sezione di Torino, Turin, Italy
57 INFN, Sezione di Trieste, Trieste, Italy
58 Inha University, Incheon, Republic of Korea
59 Institute for Gravitational and Subatomic Physics (GRASP), Utrecht University/Nikhef, Utrecht, Netherlands
60 Institute of Experimental Physics, Slovak Academy of Sciences, Košice, Slovak Republic
61 Institute of Physics, Homi Bhabha National Institute, Bhubaneswar, India
62 Institute of Physics of the Czech Academy of Sciences, Prague, Czech Republic
63 Institute of Space Science (ISS), Bucharest, Romania
64 Institut für Kernphysik, Johann Wolfgang Goethe-Universität Frankfurt, Frankfurt, Germany
65 Instituto de Ciencias Nucleares, Universidad Nacional Autónoma de México, Mexico City, Mexico
66 Instituto de Física, Universidade Federal do Rio Grande do Sul (UFRGS), Porto Alegre, Brazil
67 Instituto de Física, Universidad Nacional Autónoma de México, Mexico City, Mexico
68 iThemba LABS, National Research Foundation, Somerset West, South Africa
69 Jeonbuk National University, Jeonju, Republic of Korea
70 Johann-Wolfgang-Goethe Universität Frankfurt Institut für Informatik, Fachbereich Informatik und Mathematik, Frankfurt, Germany
71 Korea Institute of Science and Technology Information, Daejeon, Republic of Korea
72 Laboratoire de Physique Subatomique et de Cosmologie, Université Grenoble-Alpes, CNRS-IN2P3, Grenoble, France
73 Lawrence Berkeley National Laboratory, Berkeley, California, United States
74 Lund University Department of Physics, Division of Particle Physics, Lund, Sweden
75 Nagasaki Institute of Applied Science, Nagasaki, Japan
76 Nara Women’s University (NWU), Nara, Japan
77 National and Kapodistrian University of Athens, School of Science, Department of Physics , Athens, Greece
78 National Centre for Nuclear Research, Warsaw, Poland
79 National Institute of Science Education and Research, Homi Bhabha National Institute, Jatni, India
80 National Nuclear Research Center, Baku, Azerbaijan
81 National Research and Innovation Agency - BRIN, Jakarta, Indonesia
82 Niels Bohr Institute, University of Copenhagen, Copenhagen, Denmark
83 Nikhef, National institute for subatomic physics, Amsterdam, Netherlands
84 Nuclear Physics Group, STFC Daresbury Laboratory, Daresbury, United Kingdom
85 Nuclear Physics Institute of the Czech Academy of Sciences, Husinec-Řež, Czech Republic
86 Oak Ridge National Laboratory, Oak Ridge, Tennessee, United States
87 Ohio State University, Columbus, Ohio, United States
88 Physics department, Faculty of science, University of Zagreb, Zagreb, Croatia
89 Physics Department, Panjab University, Chandigarh, India
90 Physics Department, University of Jammu, Jammu, India
91 Physics Program and International Institute for Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2), Hiroshima University, Hiroshima, Japan
92 Physikalisches Institut, Eberhard-Karls-Universität Tübingen, Tübingen, Germany
93 Physikalisches Institut, Ruprecht-Karls-Universität Heidelberg, Heidelberg, Germany
94 Physik Department, Technische Universität München, Munich, Germany
95 Politecnico di Bari and Sezione INFN, Bari, Italy
96 Research Division and ExtreMe Matter Institute EMMI, GSI Helmholtzzentrum für Schwerionenforschung GmbH, Darmstadt, Germany
97 Saga University, Saga, Japan
98 Saha Institute of Nuclear Physics, Homi Bhabha National Institute, Kolkata, India
99 School of Physics and Astronomy, University of Birmingham, Birmingham, United Kingdom
100 Sección Física, Departamento de Ciencias, Pontificia Universidad Católica del Perú, Lima, Peru
101 Stefan Meyer Institut für Subatomare Physik (SMI), Vienna, Austria
102 SUBATECH, IMT Atlantique, Nantes Université, CNRS-IN2P3, Nantes, France
103 Sungkyunkwan University, Suwon City, Republic of Korea
104 Suranaree University of Technology, Nakhon Ratchasima, Thailand
105 Technical University of Košice, Košice, Slovak Republic
106 The Henryk Niewodniczanski Institute of Nuclear Physics, Polish Academy of Sciences, Cracow, Poland
107 The University of Texas at Austin, Austin, Texas, United States
108 Universidad Autónoma de Sinaloa, Culiacán, Mexico
109 Universidade de São Paulo (USP), São Paulo, Brazil
110 Universidade Estadual de Campinas (UNICAMP), Campinas, Brazil
111 Universidade Federal do ABC, Santo Andre, Brazil
112 Universitatea Nationala de Stiinta si Tehnologie Politehnica Bucuresti, Bucharest, Romania
113 University of Derby, Derby, United Kingdom
114 University of Houston, Houston, Texas, United States
115 University of Jyväskylä, Jyväskylä, Finland
116 University of Kansas, Lawrence, Kansas, United States
117 University of Liverpool, Liverpool, United Kingdom
118 University of Science and Technology of China, Hefei, China
119 University of South-Eastern Norway, Kongsberg, Norway
120 University of Tennessee, Knoxville, Tennessee, United States
121 University of the Witwatersrand, Johannesburg, South Africa
122 University of Tokyo, Tokyo, Japan
123 University of Tsukuba, Tsukuba, Japan
124 Universität Münster, Institut für Kernphysik, Münster, Germany
125 Université Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
126 Université de Lyon, CNRS/IN2P3, Institut de Physique des 2 Infinis de Lyon, Lyon, France
127 Université de Strasbourg, CNRS, IPHC UMR 7178, F-67000 Strasbourg, France, Strasbourg, France
128 Université Paris-Saclay, Centre d’Etudes de Saclay (CEA), IRFU, Départment de Physique Nucléaire (DPhN), Saclay, France
129 Université Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
130 Università degli Studi di Foggia, Foggia, Italy
131 Università del Piemonte Orientale, Vercelli, Italy
132 Università di Brescia, Brescia, Italy
133 Variable Energy Cyclotron Centre, Homi Bhabha National Institute, Kolkata, India
134 Warsaw University of Technology, Warsaw, Poland
135 Wayne State University, Detroit, Michigan, United States
136 Yale University, New Haven, Connecticut, United States
137 Yildiz Technical University, Istanbul, Turkey
138 Yonsei University, Seoul, Republic of Korea
139 Affiliated with an institute formerly covered by a cooperation agreement with CERN
140 Affiliated with an international laboratory covered by a cooperation agreement with CERN.